OpenStudy (anonymous):

4x − y = 8 2x + 3y = 9 8 − 4x 8 + 4x −8 + 4x −8 − 4x

OpenStudy (abb0t):

What's the question?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!

OpenStudy (anonymous):

For the following system, if you isolated y in the first equation to use the Substitution Method, what expression would you substitute into the second equation?

OpenStudy (anonymous):

if you solve 1st eq for y, then y=4x-8

OpenStudy (anonymous):

whts the answer

OpenStudy (anonymous):

4x − y = 8 2x + 3y = 9 ------------ multiply the top equation by 3 12x-3y = 24 2 x+3y = 9 ----------- add vertically (12+2)x +0y=24+9 14x = 33 x = 33/14 plug back into either equation to get y.